# CaRiNG: Learning Temporal Causal Representation under Non-Invertible Generation Process

### v1

@@ 4 Approach
Given our results on identifiability, we introduce our CaRiNG approach. This aims to estimate the latent causal dynamics presented in Eq 1, even when faced with a non-invertible mixing procedure. To achieve this, CaRiNG builds upon the Sequential Variational Auto-Encoders (Sequential VAE (Li & Mandt, 2018)) and incorporates three primary modules: the sequence-to-step encoder (SeqEnc), the step-to-step decoder (StepDec), and the transition prior module ($\hat{\mathbf{f}}^{-1}_{\mathbf{z}}$). Through Sequential VAE, we ensure the reconstruction capability from latent variables to observed variables. Meanwhile, in contrast to the Gaussian prior in VAEs, our method employs normalizing flow to control the prior distribution, ensuring that the latent variables satisfy the assumed conditional independence. During the training phase, we integrate the conditions from Sec. 3 as constraints and adopt two corresponding loss functions.
@@
Transition Prior Module. To uphold the conditional independence assumption, we propose to minimize the KL divergence between the posterior distribution, $q(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t:t-\mu})$, and a prior distribution $p(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-\tau})$. By hard-coding the prior distribution to be conditionally independent on $\mathbf{\hat{z}}_{t-1:t-\tau}$, we expect the posterior to also be subject to the conditional independence assumption. Direct estimation of the prior, which has an arbitrary density function, poses challenges. As a solution, we introduce a transition prior module that facilitates the estimation of the prior using normalizing flow. Specifically, the prior is represented through a Gaussian distribution combined with the Jacobian matrix of the transition module.
@@
The transition prior module can be efficiently executed using a MLP, transforming the latent variables $\mathbf{\hat{z}}_{t:t-\tau}$ into $\mathbf{\hat{\epsilon}}_{t}$.
@@
Optimization. We train CaRiNG using the Evidence Lower BOund (ELBO) objective, which is written as
$\small\begin{split}&\textsc{ELBO}\!\triangleq\!\mathbb{E}_{q_{\phi}(\mathbf{Z}|\mathbf{X})}[\log p_{\theta}(\mathbf{X}|\mathbf{Z})]-D_{KL}(q_{\phi}(\mathbf{Z}|\mathbf{X})||p_{\theta}(\mathbf{Z}))\\ &\!=\!\underbrace{\mathbb{E}_{q_{\phi}(\mathbf{Z}|\mathbf{X})}\sum_{t=1}^{T}\log p_{\theta}(\mathbf{x}_{t}|\mathbf{z}_{t})}_{-\mathcal{L}_{\text{Recon}}}\\ &\!+\!\underbrace{\mathbb{E}_{q_{\phi}(\mathbf{Z}|\mathbf{X})}\left[\sum_{t=1}^{T}\log p_{\theta}(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-\tau})\!-\!\sum_{t=1}^{T}\log q_{\phi}(\mathbf{z}_{t}|\mathbf{x}_{t:t-\mu})\right]}_{-\mathcal{L}_{\text{KLD}}}.\end{split}\vspace{-0.2cm}$ (9)
@@ 5.1 Simulation Experiments
Dataset and experimental settings. To evaluate whether CaRiNG can learn the causal process and identify the latent variables under a non-invertible scenario, we design a series of simulation experiments based on a random causal structure with a given sample size and variable size. We provide two experimental settings, including NG and NG-TDMP, which simulate the scenarios in Theorem 1 with $r=0$ (non-invertible generation) and $r>0$ (time-delayed mixing process), respectively. In particular, for NG, we simulate the visual perception system that uses the ground-truth dimension as 3 to represent the 3D real world and apply 2 measured variables to represent the 2D observation, which indicates the generation is non-invertible. For NG-TDMP, we simulate the persistence of vision that involves the previous latent variables in the current mixing process. It denotes that even if the dimension of the observation is not reduced, the generation process is still non-invertible due to the time-delay mixing. More details of the data generation process can be found in the Appendix.
@@
Table 2: Results on SUTD-TrafficQA dataset on different question types. B: “Basic understanding”, F: “Forecasting task”, R: “Reverse Reasoning”, C: “Counterfactual inference”, I: “Introspection”, A: “Attribution”. The "ALL" column shows the weighted average of scores, based on each question type’s population. The cross-modality matching parts of TDRL and CaRiNG are based on HCRN.
Method Year Question Type All
B A I C F R
I3D+LSTM CVPR2017 - - - - - - 33.21
HCRN CVPR2020 34.17 50.29 33.40 40.73 44.58 50.09 36.26
VQAC ICCV2021 34.02 49.43 34.44 39.74 38.55 49.73 36.00
MASN ACL2021 33.83 50.86 34.23 41.06 41.57 50.80 36.03
DualVGR TMM2021 33.91 50.57 33.40 41.39 41.57 50.62 36.07
Eclipse CVPR2021 - - - - - - 37.05
CMCIR TPAMI2023 36.10 52.59 38.38 46.03 48.80 52.21 38.58
TDRL NeurIPS2022 36.28 39.57 29.63 46.49 31.84 39.46 37.32
CaRiNG - 38.95 44.98 32.43 48.64 41.70 47.16 41.22
@@ 5.2 Real-world Experiments
Quantitative results. Performance comparisons for the six question types on SUTD-TrafficQA are summarized in Table 2. CaRiNG achieves a score of 41.22, which demonstrates a significant improvement which is nearly 6.8% over the next best method. Notably, when compared to TDRL, which lacks temporal context, CaRiNG exhibits significant advancements in representing complex, non-invertible traffic events. When benchmarked against the HCRN baseline, which employs the same cross-modality matching module, our approach further escalates the score by 4.96 through causal representation learning. Though CMCIR (Liu et al., 2023) applies the Swin-Transformer-L (Liu et al., 2021) pretrained on ImageNet-22K dataset as the frame-level appearance extractor and employs the video Swin-B (Liu et al., 2022) pretrained on Kinetics-600 as the clip-level motion feature extractor (more powerful than ours), CaRiNG with sample ResNet101 (He et al., 2016) features still outperforms it with 2.64 in average. More analysis on visualization, complexity, and parameters ($\tau$ and $\mu$) are included in Appendix C,  D.1,  D.3, and  D.4, respectively.
@@ 6 Conclusion
In this paper, we have proposed to consider learning temporal causal representation under the non-invertible generation process, which motivated by the common requirement of the temporal system, such as the visual perception process. We have established identifiability theories that allow for recovering the latent causal process with the nonlinear and non-invertible mixing function. Furthermore, based on this theorem, we proposed our approach, CaRiNG, to leverage the temporal context to estimate the lost latent information. We have conducted a series of simulated experiments to verify the identifiability results of CaRiNG under the non-invertible generations, and evaluated the learned representation in a complex and non-invertible traffic environment with real-world VideoQA tasks.
@@
CaRiNG: Learning Temporal Causal Representation under Non-Invertible Generation Process”
Appendix organization:
[…]
1
Appendix A Identifiability Theory
A.1 Proof for Theorem 1
@@
$\mathbf{x}_{t}=\mathbf{g}(\mathbf{z}_{t:t-r}),\quad z_{it}=f_{i}\left(\mathbf{z}_{t-1},\epsilon_{it}\right),\quad\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu}).$ (10)
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ whose probability density function is smooth and positive, suppose there exists function $\mathbf{\hat{g}},\mathbf{\hat{m}}$ which subject to observational equivalence, i.e.,
@@
[…]
    for each value of $\mathbf{z}_{t}$, $\mathbf{v}_{1t},\mathring{\mathbf{v}}_{1t},\mathbf{v}_{2t},\mathring{\mathbf{v}}_{2t},...,\mathbf{v}_{nt},\mathring{\mathbf{v}}_{nt}\in\mathbf{R}^{2n}$, as $2n$ vector functions in $z_{1,t-1}$, $z_{2,t-1}$, …, $z_{n,t-1}$, are linearly independent,
@@
For a better understanding of the sufficiency assumption in Eq 12, we will now proceed to provide an explanation for it. Consider the mixed derivative matrix $\mathbb{V}_{t},\mathbb{\mathring{V}}_{t}$ formed by $\{\mathbf{v}_{k,t}\}_{k=1}^{n}$ , $\{\mathbf{\mathring{v}}_{k,t}\}_{k=1}^{n}$ as
$\mathbb{V}_{t}=\begin{bmatrix}\frac{\partial^{2}\eta_{1t}}{\partial z_{1,t}\partial z_{1,t-r-1}}&\frac{\partial^{2}\eta_{1t}}{\partial z_{1,t}\partial z_{2,t-r-1}}&\cdots&\frac{\partial^{2}\eta_{1t}}{\partial z_{1,t}\partial z_{n,t-r-1}}\\ \frac{\partial^{2}\eta_{2t}}{\partial z_{2,t}\partial z_{1,t-r-1}}&\frac{\partial^{2}\eta_{2t}}{\partial z_{2,t}\partial z_{2,t-r-1}}&\cdots&\frac{\partial^{2}\eta_{2t}}{\partial z_{2,t}\partial z_{n,t-r-1}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial^{2}\eta_{nt}}{\partial z_{n,t}\partial z_{1,t-r-1}}&\frac{\partial^{2}\eta_{nt}}{\partial z_{n,t}\partial z_{2,t-r-1}}&\cdots&\frac{\partial^{2}\eta_{nt}}{\partial z_{n,t}\partial z_{n,t-r-1}}\\ \end{bmatrix}^{T}$ (13)
and
$\mathbb{\mathring{V}}_{t}=\begin{bmatrix}\frac{\partial^{3}\eta_{1t}}{\partial^{2}z_{1,t}\partial z_{1,t-r-1}}&\frac{\partial^{3}\eta_{1t}}{\partial^{2}z_{1,t}\partial z_{2,t-r-1}}&\cdots&\frac{\partial^{3}\eta_{1t}}{\partial^{2}z_{1,t}\partial z_{n,t-r-1}}\\ \frac{\partial^{3}\eta_{2t}}{\partial^{2}z_{2,t}\partial z_{1,t-r-1}}&\frac{\partial^{3}\eta_{2t}}{\partial^{2}z_{2,t}\partial z_{2,t-r-1}}&\cdots&\frac{\partial^{3}\eta_{2t}}{\partial^{2}z_{2,t}\partial z_{n,t-r-1}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial^{3}\eta_{nt}}{\partial^{2}z_{n,t}\partial z_{1,t-r-1}}&\frac{\partial^{3}\eta_{nt}}{\partial^{2}z_{n,t}\partial z_{2,t-r-1}}&\cdots&\frac{\partial^{3}\eta_{nt}}{\partial^{2}z_{n,t}\partial z_{n,t-r-1}}\\ \end{bmatrix}^{T}$ (14)
separately. The sufficiency assumption is satisfied if and only if $\mathbb{V}_{t},\mathbb{\mathring{V}}_{t}$ are of full rank, thus $\begin{bmatrix}\mathbb{V}_{t}&\mathbf{0}\\ \mathbf{0}&\mathbb{\mathring{V}}_{t}\end{bmatrix}$ is of full rank. The purpose of this assumption is that the model can capture and distinguish independent noise from transition dynamics only when they are sufficiently diverse. This property will be used in the following proof.
@@
For any $t$, combining Eq 10 and Eq 11 gives
$\mathbf{z}_{t}$ $=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ (15)
@@
as well as $\mathbf{\hat{z}}_{t}=\mathbf{\hat{m}}(\mathbf{g}(\mathbf{z}_{t},\mathbf{m}(\mathbf{x}_{t-1:t-\mu-1}),\cdots,\mathbf{m}(\mathbf{x}_{t-r:t-\mu-r})),\mathbf{x}_{t-1:t-\mu})$ similarly. Upon Eq 15, we have an unified partially invertible function $\mathbf{z}_{t}=\mathbf{h}(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t-1:t-\mu-r})$ where $\mathbf{h}=\mathbf{m}\circ\hat{\mathbf{g}}$ with Jacobian $\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}=\mathbf{H}_{t}(\mathbf{\hat{z}}_{t};\mathbf{x}_{t-1:t-\mu-r})$. By partially invertible it means that $\mathbf{z}$ and $\mathbf{\hat{z}}$ are in one-to-one correspondence for any context observations $\mathbf{x}_{t-1:t-\mu-r}$ that are fixed. One more thing to notify is that since $\mathbf{g},\mathbf{\hat{g}},\mathbf{m},\mathbf{\hat{m}}$ are second-order differentiable, the nested $\mathbf{h}$ is also second-order differentiable. Let us consider the mapping from joint distribution $(\mathbf{\hat{z}}_{t},\mathbf{x}_{t-1:t-\mu-r-1})$ to $(\mathbf{z}_{t},\mathbf{x}_{t-1:t-\mu-r-1})$, i.e.,
$P(\mathbf{z}_{t},\mathbf{x}_{t-1:t-\mu-r-1})=P(\mathbf{\hat{z}}_{t},\mathbf{x}_{t-1:t-\mu-r-1})\,/\,|\mathbf{J}_{t}|,$ (16)
@@
$\mathbf{J}_{t}=\begin{bmatrix}\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}&\mathbf{0}\\ *&\mathbf{I}\end{bmatrix},$ (17)
which is a lower triangle matrix, where $\mathbf{I}$ infers eye matrix and *** infers any possible matrix. Thus, we have determinant $|\mathbf{J}_{t}|=|\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}|=|\mathbf{H}_{t}|$. Dividing both sides of Eq 16 by $P(\mathbf{x}_{t-1:t-\mu-r-1})$ gives
$\textbf{LHS}=P(\mathbf{z}_{t}|\mathbf{x}_{t-1:t-\mu-r-1})=P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-1}),$ (18)
@@
$P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-1})=P(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-r-1})\,/\,|\mathbf{H}_{t}|.$ (19)
@@
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}=0.$ (20)
Referencing Eq 19, it gets expressed as:
$\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1})=\log p({\mathbf{z}}_{t}\,|\,{\mathbf{z}}_{t-1:t-r-1})+\log|\mathbf{H}_{t}|=\sum_{k=1}^{n}\eta_{kt}+\log|\mathbf{H}_{t}|.$ (21)
@@
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}$ $=\sum_{k=1}^{n}\Big{(}\frac{\partial^{2}\eta_{kt}}{\partial z_{kt}^{2}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial\eta_{kt}}{\partial z_{kt}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}+\frac{\partial^{2}\log|\mathbf{H}_{t}|}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}.$ (22)
@@
According to Eq 20, the right-hand side of the presented equation consistently equals 0. Therefore, for each index $l$ ranging from 1 to $n$, and every associated value of $z_{l,t-r-1}$, its partial derivative with respect to $z_{l,t-r-1}$ remains 0. That is,
$\sum_{k=1}^{n}\Big{(}\frac{\partial^{3}\eta_{kt}}{\partial z_{kt}^{2}\partial z_{l,t-r-1}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial^{2}\eta_{kt}}{\partial z_{kt}\partial z_{l,t-r-1}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}\equiv 0,$ (23)
@@
where we leveraged the fact that entries of $\mathbf{H}_{t}$ do not depend on $z_{l,t-r-1}$. Considering any given value of $\mathbf{z}_{t}$, $\mathbf{v}_{1t},\mathring{\mathbf{v}}_{1t},\mathbf{v}_{2t},\mathring{\mathbf{v}}_{2t},...,\mathbf{v}_{nt},\mathring{\mathbf{v}}_{nt}$ are linearly independent, to make the above equation hold true, one has to set $\mathbf{H}_{kit}\mathbf{H}_{kjt}=0$ or $i\neq j$. In other words, each row of $\mathbf{H}_{t}$ consists of at most a single non-zero entry, and $\mathbf{z}_{t}$ must be a component-wise transformation of a permuted version of $\mathbf{\hat{z}}_{t}$. ∎

### v2

@@ 4 Approach
Given our results on identifiability, we introduce our CaRiNG approach. This aims to estimate the latent causal dynamics presented in Eq 1, even when faced with a non-invertible mixing procedure. To achieve this, CaRiNG builds upon the Sequential Variational Auto-Encoders (Sequential VAE (Chung et al., 2015; Li & Mandt, 2018)) and incorporates three primary modules: the sequence-to-step encoder (SeqEnc), the step-to-step decoder (StepDec), and the transition prior module ($\hat{\mathbf{f}}^{-1}_{\mathbf{z}}$). Through Sequential VAE, we ensure the reconstruction capability from latent variables to observed variables. Meanwhile, in contrast to the Gaussian prior in VAEs, our method employs normalizing flow to control the prior distribution, ensuring that the latent variables satisfy the assumed conditional independence. During the training phase, we integrate the conditions from Sec. 3 as constraints and adopt two corresponding loss functions.
@@
Transition Prior Module. To uphold the conditional independence assumption, we propose to minimize the KL divergence between the posterior distribution and a hard-coding prior distribution with such property. The constraint indicates that current latent variables are mutually independent, conditioned on historical latent variables. Formally, by hard-coding the prior distribution we enforce $\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-\tau}$ to be mutually independent. By minimizing the KL divergence, we expect the posterior to be subject to the assumption as well, such as $\mathbf{\hat{z}}_{t}|\mathbf{\hat{x}}_{t:t-\mu},\mathbf{\hat{z}}_{t-1:t-\tau}$ are mutually independent. Direct estimation of the prior, which has an arbitrary density function, poses challenges. As a solution, we introduce a transition prior module that facilitates the estimation of the prior using normalizing flow. Specifically, the prior is represented through a Gaussian distribution combined with the Jacobian matrix of the transition module.
@@
The transition prior module can be efficiently executed using an MLP, transforming the latent variables $\mathbf{\hat{z}}_{t:t-\tau}$ into $\mathbf{\hat{\epsilon}}_{t}$.
@@
Optimization. We train CaRiNG using the Evidence Lower BOund (ELBO) objective, which is written as follows:
$\small\begin{split}&\textsc{ELBO}\!\triangleq\!\mathbb{E}_{q_{\phi}(\mathbf{Z}|\mathbf{X})}[\log p_{\theta}(\mathbf{X}|\mathbf{Z})]-D_{KL}(q_{\phi}(\mathbf{Z}|\mathbf{X})||p_{\theta}(\mathbf{Z}))\\ &\!=\!\underbrace{\mathbb{E}_{q_{\phi}(\mathbf{Z}|\mathbf{X})}\sum_{t=1}^{T}\log p_{\theta}(\mathbf{x}_{t}|\mathbf{z}_{t})}_{-\mathcal{L}_{\text{Recon}}}\\ &\!+\!\underbrace{\mathbb{E}_{q_{\phi}(\mathbf{Z}|\mathbf{X})}\left[\sum_{t=1}^{T}\log p_{\theta}(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-\tau})\!-\!\sum_{t=1}^{T}\log q_{\phi}(\mathbf{z}_{t}|\mathbf{x}_{t:t-\mu})\right]}_{-\mathcal{L}_{\text{KLD}}}.\end{split}$ (9)
@@ 5.1 Simulation Experiments
Dataset and experimental settings. To evaluate whether CaRiNG can learn the causal process and identify the latent variables under a non-invertible scenario, we design a series of simulation experiments based on a random causal structure with a given sample size and variable size. We provide two experimental settings, including NG and NG-TDMP, which simulate the scenarios in Theorem 1 with $r=0$ (non-invertible generation) and $r>0$ (time-delayed mixing process), respectively. In particular, for NG, we simulate the visual perception system that uses the ground-truth dimension as 3 to represent the 3D real world and apply 2 measured variables to represent the 2D observation, which indicates the generation is non-invertible. For NG-TDMP, we simulate the persistence of vision that involves the previous latent variables in the current mixing process. It denotes that even if the dimension of the observation is not reduced, the generation process is still non-invertible due to the time-delay mixing. More details of the data generation process can be found in Appendix A2.1.
@@
Table 2: Results on the SUTD-TrafficQA dataset. The cross-modality matching parts of TDRL and CaRiNG are based on HCRN.
[…]
I3D+LSTM CVPR2017 33.21
HCRN CVPR2020 36.26
VQAC ICCV2021 36.00
MASN ACL2021 36.03
DualVGR TMM2021 36.07
Eclipse CVPR2021 37.05
CMCIR TPAMI2023 38.58
TDRL NeurIPS2022 37.32
CaRiNG - 41.22
@@ 5.2 Real-world Experiments
Quantitative results. Performance comparisons for the six question types on SUTD-TrafficQA are summarized in Table 2. CaRiNG achieves a score of 41.22, which demonstrates a significant improvement which is nearly 6.8% over the next best method. Notably, when compared to TDRL, which lacks temporal context, CaRiNG exhibits significant advancements in representing complex, non-invertible traffic events. When benchmarked against the HCRN baseline, which employs the same cross-modality matching module, our approach further escalates the score by 4.96 through causal representation learning. Though CMCIR (Liu et al., 2023) applies the Swin-Transformer-L (Liu et al., 2021) pretrained on ImageNet-22K dataset as the frame-level appearance extractor and employs the video Swin-B (Liu et al., 2022) pretrained on Kinetics-600 as the clip-level motion feature extractor (more powerful than ours), CaRiNG with sample ResNet101 (He et al., 2016) features still outperforms it with 2.64 in average. More analysis on TrafficQA and another evaluation on Volleyball (Ibrahim et al., 2016) can be found in Appendix A3 and A4, respectively.
@@ 6 Conclusion
In this paper, we have proposed to consider learning temporal causal representation under the non-invertible generation process, which is motivated by the common requirement of the temporal system, such as the visual perception process. We have established identifiability theories that allow for recovering the latent causal process with the nonlinear and non-invertible mixing function. Furthermore, based on this theorem, we proposed our approach, CaRiNG, to leverage the temporal context to estimate the lost latent information. We have conducted a series of simulated experiments to verify the identifiability results of CaRiNG under the non-invertible generations and evaluated the learned representation in a complex and non-invertible traffic environment with real-world VideoQA tasks.
Impact Statement
Acknowledgments
We would like to acknowledge the support from NSF Grant 2229881, the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Apple Inc., KDDI Research Inc., Quris AI, and Florin Court Capital.
@@ Acknowledgments
“Learning Temporal Causal Representation under Non-Invertible Generation Process”
Appendix A1 Identifiability Theory
A1.1 Proof for Theorem 1
@@
$\mathbf{x}_{t}=\mathbf{g}(\mathbf{z}_{t:t-r}),\quad z_{it}=f_{i}\left(\mathbf{z}_{t-1:t-r-1},\epsilon_{it}\right),\quad\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu}).$ (A1)
Theorem A1 (Identifiability under Non-invertible Generative Process).
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{d}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ , suppose there exists function $\mathbf{\hat{g}},\mathbf{\hat{m}}$ which is subject to observational equivalence,
$\mathbf{x}_{t}=\mathbf{\hat{g}}(\mathbf{\hat{z}}_{t:t-r}),\quad\mathbf{\hat{z}}_{t}=\mathbf{\hat{m}}(\mathbf{x}_{t:t-\mu}).$ (A2)
@@
  • {(Smooth and Positive Density) the probability density function of latent variables is third-order differentiable and positive in $R^{n}$,
@@
    $\mathbf{v}_{lt}\triangleq\Big{(}\frac{\partial^{2}\eta_{1t}}{\partial z_{1t}\partial z_{l,t-r-1}},...,\frac{\partial^{2}\eta_{nt}}{\partial z_{nt}\partial z_{l,t-r-1}},\frac{\partial^{3}\eta_{1t}}{\partial z_{1t}^{2}\partial z_{l,t-r-1}},...,\frac{\partial^{3}\eta_{nt}}{\partial z_{nt}^{2}\partial z_{l,t-r-1}}\Big{)}^{\intercal},$ (A3)
    for $l=1,2,\cdots,n$. For each value of $\mathbf{z}_{t}$, there exists $2n$ different values of $z_{l,t-r-1}$ such that the $2n$ vector functions $\mathbf{v}_{lt}\in\mathbf{R}^{2n}$ are linearly independent,
@@
For any $t$, combining Eq A1 and Eq A2 gives
$\mathbf{z}_{t}$ $=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ (A4)
@@
as well as $\mathbf{\hat{z}}_{t}=\mathbf{\hat{m}}(\mathbf{g}(\mathbf{z}_{t},\mathbf{m}(\mathbf{x}_{t-1:t-\mu-1}),\cdots,\mathbf{m}(\mathbf{x}_{t-r:t-\mu-r})),\mathbf{x}_{t-1:t-\mu})$ similarly. Upon Eq A4, we have an unified partially invertible function $\mathbf{z}_{t}=\mathbf{h}(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t-1:t-\mu-r})$ where $\mathbf{h}=\mathbf{m}\circ\hat{\mathbf{g}}$ with Jacobian $\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}=\mathbf{H}_{t}(\mathbf{\hat{z}}_{t};\mathbf{x}_{t-1:t-\mu-r})$. By partially invertible it means that $\mathbf{z}$ and $\mathbf{\hat{z}}$ are in one-to-one correspondence for any context observations $\mathbf{x}_{t-1:t-\mu-r}$ that are fixed. One more thing to notify is that since $\mathbf{g},\mathbf{\hat{g}},\mathbf{m},\mathbf{\hat{m}}$ are second-order differentiable, the nested $\mathbf{h}$ is also second-order differentiable. Let us consider the mapping from joint distribution $(\mathbf{\hat{z}}_{t},\mathbf{x}_{t-1:t-\mu-r-1})$ to $(\mathbf{z}_{t},\mathbf{x}_{t-1:t-\mu-r-1})$, i.e.,
$P(\mathbf{z}_{t},\mathbf{x}_{t-1:t-\mu-r-1})=P(\mathbf{\hat{z}}_{t},\mathbf{x}_{t-1:t-\mu-r-1})\,/\,|\mathbf{J}_{t}|,$ (A5)
@@
$\mathbf{J}_{t}=\begin{bmatrix}\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}&\mathbf{0}\\ *&\mathbf{I}\end{bmatrix},$ (A6)
which is a lower triangle matrix, where $\mathbf{I}$ infers eye matrix and * infers any possible matrix. Thus, we have determinant $|\mathbf{J}_{t}|=|\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}|=|\mathbf{H}_{t}|$. Dividing both sides of Eq A5 by $P(\mathbf{x}_{t-1:t-\mu-r-1})$ gives
$\textbf{LHS}=P(\mathbf{z}_{t}|\mathbf{x}_{t-1:t-\mu-r-1})=P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-1}),$ (A7)
@@
$P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-1})=P(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-r-1})\,/\,|\mathbf{H}_{t}|.$ (A8)
@@
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}=0.$ (A9)
Referencing Eq A8, it gets expressed as:
$\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1})=\log p({\mathbf{z}}_{t}\,|\,{\mathbf{z}}_{t-1:t-r-1})+\log|\mathbf{H}_{t}|=\sum_{k=1}^{n}\eta_{kt}+\log|\mathbf{H}_{t}|.$ (A10)
@@
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}$ $=\sum_{k=1}^{n}\Big{(}\frac{\partial^{2}\eta_{kt}}{\partial z_{kt}^{2}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial\eta_{kt}}{\partial z_{kt}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}+\frac{\partial^{2}\log|\mathbf{H}_{t}|}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}.$ (A11)
@@
According to Eq A9, the right-hand side of the presented equation consistently equals 0. Therefore, for each index $l$ ranging from 1 to $n$, and every associated value of $z_{l,t-r-1}$, its partial derivative with respect to $z_{l,t-r-1}$ remains 0. That is,
$\sum_{k=1}^{n}\Big{(}\frac{\partial^{3}\eta_{kt}}{\partial z_{kt}^{2}\partial z_{l,t-r-1}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial^{2}\eta_{kt}}{\partial z_{kt}\partial z_{l,t-r-1}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}\equiv 0,$ (A12)
@@
where we leveraged the fact that entries of $\mathbf{H}_{t}$ do not depend on $z_{l,t-r-1}$. Considering any given value of $\mathbf{z}_{t}$, there exists at least $2n$ different values of $\mathbf{v}_{lt}$ such that they are linearly independent. To make the above equation hold true, one has to set $\mathbf{H}_{kit}\mathbf{H}_{kjt}=0$ or $i\neq j$. In other words, each row of $\mathbf{H}_{t}$ consists of at most a single non-zero entry, and $\mathbf{z}_{t}$ must be a component-wise transformation of a permuted version of $\mathbf{\hat{z}}_{t}$. ∎